OpenStudy (anonymous):

What is the 20th term of the sequence that begins −3, 6, −12, 24 ...?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

It is a GP...... Here a=-3,common ratio r=-2 There fore using the eqn., 20th term of the sequence =a(r)^(n-1)=-3 * (-2)^19=-1572864

OpenStudy (mertsj):

Just keep multiplying y -2 until you get the 20th term. That's the easy way. Or you could use your formula.

OpenStudy (anonymous):

Got it!!!!

OpenStudy (anonymous):

@xkylex

OpenStudy (mertsj):

\[a _{20}=a _{1}r ^{20-1}\]

OpenStudy (anonymous):

Mertsji, using your formula, would I get the same result as ashwinjohn3?

OpenStudy (mertsj):

yes. He used the same formula

OpenStudy (anonymous):

Ok. Thanks all of you!

OpenStudy (mertsj):

yw